Two cards are drawn from a well-shuffled deck of 52 playing cards. Let X denote the number of aces drawn. Find P(X = 2). (Round your answer to three decimal places.)
Answer by stanbon(57214)

Two cards are drawn from a well-shuffled deck of 52 playing cards. Let X denote the number of aces drawn. Find P(X = 2).
----------------------
Number of ways to succeed: 4C2 = 6
---
Number of random pairs: 52C2 = (52*51)/(1*2) = 26*51 =
----
P(x = 2 aces) = 6/52C2 = 6/1326 = 3/663 = 1/221
---------------------------------------------------
Cheers,
Stan H.